OpenStudy (anonymous):

Ruby was doing a growth experiment for school and found that her hair was growing at a rate of 3/20 centimeters per week. How many centimeters will her hair grow in 1/3 of a week? 9/60 20/9 9/20 1/20

OpenStudy (anonymous):

ok first we have to sort out what is wat the rate = 3/20 the time is = 1/3 so in order to get the total length, we just multiply how fast her hair grows/per week, by the number of weeks: 3/20 cm per week x 1/3 week = 3/60 cm = 1/20 cm